# quality of service 5G fifth generation

### v2

@@ III System Model and Problem Formulation
In our study, we explore a deployment based on O-RAN, wherein each Base Station (BS), known as RU in O-RAN, supports two different slices: eMBB and URLLC. These slices are designed to meet the QoS requirements of User Equipments. Our analysis focuses on two key network functions, power control, and radio resource allocation, which are jointly considered. We represent the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$. Each BS $b$ serves two sets of UEs: $\mathcal{U}_{e,b}$ for eMBB and $\mathcal{U}_{u,b}$ for URLLC, leading to a total of $U$ UEs, where $U=U_{e,b}+U_{u,b}$. Each UE is exclusively associated with a single RU. The temporal and spectral layout is divided into $\mathcal{M}$ segments known as PRBs, which are the smallest unit of resources for allocation [32]. Each PRB covers a specific time slot in the temporal domain and a frequency range of f Hz in the spectral domain. We denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices. The total number of slices ($S$) will be $S=S_{e}+S_{u}$.
The UE association is indicated by a binary variable $\alpha_{u,b,s}$, which takes a value of 1 if UE $u$ is connected to RU $b$ in slice $s$, and 0 if not. Since each UE should be connected to only one RU, the following constraint must be met:
$\sum_{b\in\mathcal{B}}\alpha_{u,b,s}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (1)
We define the binary variable $\beta_{u,b,m,s}$ as the indicator for PRB allocation, taking the value of 1 when UE $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise.
$\beta_{u,b,m,s}\leq\alpha_{u,b,s}\;\;\;\;\forall b\in\mathcal{B},\;\;m\in\mathcal{M},\;\;u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (2)
Similarly, each PRB $m$ of RU $b$ should be associated with only one UE, as defined below:
$\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},\;\;s\in\mathcal{S}.$ (3)
We denote the transmission power of PRB $m$ in RU $b$ in slice $s$ as $p_{u,b,m,s}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u,b,m,s}$. The rate achieved for a UE $u$ in slice $s$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u,b,m,s}$, where $\eta_{u,b,m,s}$ represents the Signal to Interference & Noise Ratio (SINR). The calculation for $R_{u,b,m,s}$ is determined by:
$\eta_{u,b,m,s}=\frac{\alpha_{u,b,s}h_{u,b,m,s}p_{u,b,m,s}}{\sum_{j\neq b}^{N}\sum_{l\neq s}^{S}\sum_{i\neq u}^{U}\alpha_{i,j,l}h_{i,j,m,l}p_{i,j,m,l}+\sigma^{2}},$ (4)
$R_{u,b,m,s}=\log_{2}(1+\eta_{u,b,m,s}),$ (5)
$R_{u,s}=\sum_{b\in B}\sum_{m\in M}\alpha_{u,b,s}\beta_{u,b,m,s}R_{u,b,m,s},$ (6)
$R_{u}$ $=\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}\alpha_{u,b,m,s}\sum_{m\in\mathcal{M}}\beta_{u,b,m,s}R_{u,b,m,s}$
$=\sum_{u\in\mathcal{U}}\sum_{m\in\mathcal{M}}\beta_{u,b,m}R_{u,b,m,s}\;\;\;\;\forall m\in\mathcal{M},\;s\in\mathcal{S}.$ (7)
The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P_{b}$ and $C_{b}$, respectively. Quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, mainly when the original signal displays significant variations or spans a wide range of values. This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE. These values are determined as follows:
$P_{b}=\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u,b,s}h_{u,b,m,s}p_{u,b,m,s}}+\sigma_{1}^{2},$ (8)
$C_{b}=\log_{2}(1+\frac{\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u,b,s}h_{u,b,m,s}p_{b,m,s}}}{\sigma_{1}^{2}})$
$=\log_{2}(\frac{P_{b}}{\sigma_{1}^{2}}),$ (9)
The channel gain matrices in our model are generated using a Rayleigh fading process to simulate variable channel conditions typical in non-line-of-sight environments. The channel vector from the PRB $m$ in RU $b$ to the $u$ UE in the $s$ slice, denoted by $\mathbf{h}_{u,b,m,s}$, is modeled as $\mathbf{h}_{u,b,m,s}=\sqrt{\beta_{u,b,m,s}}\mathbf{g}_{u,b,m,s}$, where $\beta_{u,b,m,s}$ represents the large-scale fading coefficient, accounting for path loss and shadowing, and $\mathbf{g}_{u,b,m,s}\sim\mathcal{N}(0,N_{0}I_{D_{s}})$ is the fast and flat fading effects modeled as a circularly symmetric complex Gaussian random variable. To reflect the impact of interference and uncertainties, the channel gain matrix $\mathbf{H}$ is augmented with Gaussian noise, expressed as $\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and adjustable variance $\sigma^{2}$.
The total delay experienced by each UE is the sum of propagation and transmission delays. The overall propagation delay ($D^{pro}$) consists of the delays in the fronthaul ($D^{fr,p}$), midhaul ($D^{mid,p}$), and backhaul ($D^{b,p}$) links. In each of these links, the propagation delay is determined as the time it takes for a signal to traverse the distance, calculated as $D=\frac{L}{C}$, where $L$ is the link length and $C$ is the propagation speed. Similarly, the total transmission delay ($D^{tr}$) is the sum of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{b,t}$) links. Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, determined by $D=\frac{\mu}{R_{u,s}}$, where $R_{u,s}$ is the data rate for UE $u$ in slice $s$, and $\mu$ is the mean packet size. The total delay for UE $u$ in slice $s$ is $D_{u,s}=\frac{L}{C}+\frac{\mu}{R_{u,s}}$, and should be below a specific threshold ($D^{max}_{u,s}$), expressed as $D_{u,s}\leq D^{max}_{u,s}$.
The objective function is designed to optimize the overall network utility, which involves maximizing the weighted sum of the data rates achieved by UEs while ensuring satisfying the QoS requirements. In this context, $w_{s}$ indicates the priority associated with individual service types. The intelligent RU is responsible for effectively managing the requirements of both network slices. Hence, we can mathematically express the problem in the following manner:
$\max_{\mathbf{p,\alpha,\beta}}$ $\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}w_{s}R_{u,s}$ (10)
subject to
$P_{b}\leq P_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (11)
$p_{u,b,m,s}\geq 0\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (12)
$p_{u,b,m,s}\leq P_{s}^{max}\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (13)
$R_{u,s}\geq R_{s}^{min}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (14)
$C_{b}\leq C_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (15)
$D_{u,s}\leq D^{max}_{u,s}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S},$ (16)
$\alpha_{u,b,s}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},s\in\mathcal{S}$ (17)
$\beta_{u,b,m,s}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (18)
$\sum_{b\in\mathcal{B}}\alpha_{u,b,s}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}$ (19)
$\beta_{u,b,m,s}\leq\alpha_{u,b,s}\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},u\in\mathcal{U},s\in\mathcal{S}$ (20)
$\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (21)
The system model is constrained by limited resources such as UE or RU and slice power and fronthaul capacity, preventing the aggregate throughput from exceeding its optimal value. This ensures that the objective function, representing aggregate throughput, has an upper bound and cannot increase infinitely. As a result, the algorithm used to solve the system will converge to an optimal solution. If the objective function is a strictly ascending function concerning the number of iterations, it will converge to its global optimum. The algorithm will converge to a local optimum if it is a non-monotonically ascending function. This guarantees the stability and convergence of the system model within the feasible region.
@@
$b,\mathcal{B},B$ Index, set, and the number of RUs
@@
$\alpha_{u,b,s}$ UE $u$ association to RU $b$ in slice $s$
$\beta_{u,b,m,s}$ UE $u$ association to PRB $m$ of RU $b$ in slice $s$
$p_{u,b,m,s}$ Transmission power of PRB $m$ of RU $b$ in slice $s$
$R_{u,b,m,s}$ Rate achieved for UE $u$ in PRB $m$ of RU $b$ in slice $s$
$R_{u,s}$ Rate achieved for UE $u$ in slice $s$
@@
$P_{b}$ Transmitted power RU in the fronthaul link
$C_{b}$ Data rate in the fronthaul link
@@
$D^{pro}$ Overall propagation delay
@@
$D^{b,p}$ Propagation delay in the backhaul link
@@
$D^{b,t}$ Transmission delay in the backhaul link
@@
TABLE I: Notations with definitions for the system model parameters.
IV Methodology
Our research addresses the resource allocation problem within the O-RAN architecture by introducing a semi-supervised learning algorithm. To achieve this, we have compiled a comprehensive dataset that includes various resource allocation scenarios, covering different objective functions and constraint combinations. We have created labeled data (computed $\alpha$, $\beta$, and transmission powers) using exhaustive search algorithm (ESA), thoroughly exploring the feasible solution space to find the global optimum solution as input for our algorithm’s supervised learning part and as a benchmark for models evaluation.
We use a deep neural network (DNN) model with a contrastive loss function to solve the resource allocation problem. This DNN model effectively manages the complexities of resource allocation and acquires the ability to map input features to resource allocation decisions. By training our model, we fine-tune it to estimate resource allocation decisions that fulfill the requirements of the optimization problem (QoS). This semi-supervised learning algorithm, which combines a VAE with a contrastive loss, is denoted as ”SS-VAE.”
We aim to evaluate the performance of our SS-VAE method using multiple regression-based performance metrics to ensure its robustness. We will also validate the effectiveness of our model by comparing its results with those of the ESA across various scenarios. While ESA is an algorithm for resource allocation decisions that find global optimum solutions, its complexity makes it impractical for real-world deployment. Therefore, we will also compare its performance with state-of-the-art methods known for their practical deployment capabilities, such as DRL. This thorough evaluation will demonstrate the reliability of our method across different scenarios.
Figure 2 illustrates our proposed methodology for training the SS-VAE and comparing its results with those of the benchmark algorithms.

### v3

@@ 4 System Model
$b,\mathcal{B},B$ Index, set, and the number of Radio Units
@@
$\alpha_{u_{s}}^{b}$ UE $u$ association to RU $b$ in slice $s$
$\beta_{u_{s},m}^{b}$ UE $u$ association to PRB $m$ of RU $b$ in slice $s$
$p_{u_{s},m}^{b}$ Transmission power of PRB $m$ of RU $b$ in slice $s$
$R_{u_{s},m}^{b}$ Rate achieved for UE $u$ in PRB $m$ of RU $b$ in slice $s$
$R_{u_{s}}$ Rate achieved for UE $u$ in slice $s$
@@
$P^{b}$ Transmitted power RU in the fronthaul link
$C^{b}$ Data rate in the fronthaul link
@@
$D^{\text{pro}}$ Overall propagation delay
@@
$D^{bc,p}$ Propagation delay in the backhaul link
@@
$D^{bc,t}$ Transmission delay in the backhaul link
@@
TABLE II: Notations with definitions for the system model parameters.
In our study, we explore a deployment based on O-RAN, wherein each Base Station (BS), known as RU in O-RAN, supports two different slices: eMBB and URLLC. These slices are designed to meet the QoS requirements of UEs. Our analysis focuses on two key network functions, power control and radio resource allocation, which are jointly considered.
We represent the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$. Each RU $b$ serves two sets of UEs: $\mathcal{U}_{e}^{b}$ for eMBB and $\mathcal{U}_{u}^{b}$ for URLLC, leading to a total of $U$ UEs, where $U=U_{e}^{b}+U_{u}^{b}$. Each UE is exclusively associated with a single RU. The temporal and spectral layout is divided into $\mathcal{M}$ segments known as PRBs, which are the smallest unit of resources for allocation [39]. Each PRB covers a specific time slot in the temporal domain and a frequency range of f Hz in the spectral domain. We denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices.
The UE association is indicated by a binary variable $\alpha_{u_{s}}^{b}$ taking a value of 1 if UE $u$ in slice $s$ is connected to RU $b$ and 0 otherwise. Since each UE should be connected to only one RU, the following constraint must be met:
$\sum_{b\in\mathcal{B}}\alpha_{u_{s}}^{b}=1,\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (1)
We define the binary variable $\beta_{u_{s},m}^{b}$ as the indicator for PRB allocation, taking the value of 1 when UE $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise. We have
$\beta_{u_{s},m}^{b}\leq\alpha_{u_{s}}^{b},\;\;\;\;\forall b\in\mathcal{B},\;\;m\in\mathcal{M},\;\;u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (2)
Similarly, each PRB $m$ of RU $b$ is associated with only one UE, as
$\sum_{u\in\mathcal{U}}\alpha_{u_{s}}^{b}\beta_{u_{s},m}^{b}\leq 1,\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},\;\;s\in\mathcal{S}.$ (3)
We denote the transmission power of PRB $m$ in RU $b$ in slice $s$ as $p_{u_{s},m}^{b}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u_{s},m}^{b}$. The rate achieved for UE $u$ in slice $s$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u_{s},m}^{b}$, where $\eta_{u_{s},m}^{b}$ represents the Signal to Interference & Noise Ratio (SINR). The calculation for $R_{u_{s},m}^{b}$ is determined by
$\eta_{u_{s},m}^{b}=\frac{\alpha_{u_{s}}^{b}h_{u_{s},m}^{b}p_{u_{s},m}^{b}}{\sum_{j\neq b}^{N}\sum_{l\neq s}^{S}\sum_{i\neq u}^{U}\alpha_{i_{l}}^{j}h_{i_{l},m}^{j}p_{i_{l},m}^{j}+\sigma^{2}},$ (4)
$R_{u_{s},m}^{b}=\log_{2}(1+\eta_{u_{s},m}^{b}),$ (5)
$R_{u_{s}}=\sum_{b\in B}\sum_{m\in M}\alpha_{u_{s}}^{b}\beta_{u_{s},m}^{b}R_{u_{s},m}^{b},$ (6)
$R_{u}$ $=\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}\alpha_{u,b,m,s}\sum_{m\in\mathcal{M}}\beta_{u_{s},m}^{b}R_{u,b,m,s}$
$=\sum_{u\in\mathcal{U}}\sum_{m\in\mathcal{M}}\beta_{u,m}^{b}R_{u_{s},m}^{b}\;\;\;\;\forall m\in\mathcal{M},\;s\in\mathcal{S}.$ (7)
The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P^{b}$ and $C^{b}$, respectively. Quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, mainly when the original signal displays significant variations or spans a wide range of values. This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE. These values are determined as follows:
$P^{b}=\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u_{s}}^{b}h_{u_{s},m}^{b}p_{u_{s},m}^{b}}+\sigma_{1}^{2},$ (8)
$C^{b}=\log_{2}(1+\frac{\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u_{s}}^{b}h_{u_{s},m}^{b}p_{s,m}^{s}}}{\sigma_{1}^{2}})$
$=\log_{2}(\frac{P_{b}}{\sigma_{1}^{2}}).$ (9)
The channel characteristics in our model are simulated using a Rayleigh fading process, which represents the variable channel conditions typical of non-line-of-sight (NLOS) environments. The channel vector from PRB $m$ in RU $b$ to $u$ UE in $s$ slice, denoted by $\mathbf{h}_{u_{s},m}^{b}$, is modeled as $\mathbf{h}_{u_{s},m}^{b}=\sqrt{\beta_{u_{s},m}^{b}}\mathbf{g}_{u_{s},m}^{b}$, where $\beta_{u_{s},m}^{b}$ represents the large-scale fading coefficient, accounting for path loss and shadowing, and $\mathbf{g}_{u_{s},m}^{b}\sim\mathcal{N}(0,N_{0}I_{D_{s}})$ is the fast and flat fading effects modeled as a circularly symmetric complex Gaussian random variable. To reflect the impact of interference and uncertainties, the channel gain matrix $\mathbf{H}$ is augmented with Gaussian noise, expressed as $\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and adjustable variance $\sigma^{2}$.
The total delay experienced by each UE is the sum of propagation and transmission delays. The overall propagation delay ($D^{\text{pro}}$) consists of the delays in the fronthaul ($D^{fr,p}$), midhaul ($D^{mid,p}$), and backhaul ($D^{bc,p}$) links. In each of these links, the propagation delay is determined as the time it takes for a signal to traverse the distance, calculated as $D=L/C$, where $L$ is the link length and $C$ is the propagation speed. Similarly, the total transmission delay ($D^{tr}$) is the sum of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{bc,t}$) links. Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, determined by $D=\frac{\mu}{R_{u_{s}}}$, where $R_{u_{s}}$ is the data rate for UE $u$ in slice $s$, and $\mu$ is the mean packet size. The total delay for UE $u$ in slice $s$ is $D_{u_{s}}=L/C+\mu/R_{u_{s}}$, and should be below a specific threshold ($D^{\max}_{u_{s}}$), expressed as $D_{u_{s}}\leq D^{\max}_{u_{s}}$.
The objective of our resource allocation scheme is to optimize the overall network utility. This primarily involves maximizing the weighted sum of the data rates, $R_{u_{s}}$, achieved by all UEs, while simultaneously satisfying their respective QoS requirements, particularly the delay constraints. The weight $w_{s}$ reflects the priority or importance associated with service type $s$ (e.g., eMBB or URLLC). The intelligent resource allocation function, managed by xApps at the RIC and influencing RU behavior, is responsible for effectively managing these diverse requirements. Hence, we can mathematically express the optimization problem as follows:
$\max_{\mathbf{p,\alpha,\beta}}$ $\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}w_{s}R_{u_{s}}$ (10)
subject to
$P^{b}\leq P_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (11)
$p_{u_{s},m}^{b}\geq 0\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (12)
$p_{u_{s},m}^{b}\leq P_{s}^{max}\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (13)
$R_{u_{s}}\geq R_{s}^{min}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (14)
$C^{b}\leq C_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (15)
$D_{u_{s}}\leq D^{max}_{u_{s}}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (16)
$\alpha_{u_{s}}^{b}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},s\in\mathcal{S}$ (17)
$\beta_{u_{s},m}^{b}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (18)
$\sum_{b\in\mathcal{B}}\alpha_{u_{s}}^{b}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}$ (19)
$\beta_{u_{s},m}^{b}\leq\alpha_{u_{s}}^{b}\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},u\in\mathcal{U},s\in\mathcal{S}$ (20)
$\sum_{u\in\mathcal{U}}\alpha_{u_{s}}^{b}\beta_{u_{s},m}^{b}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (21)
Equation (10) defines the objective of maximizing the total weighted data rate across all users and services. Constraints (11), (12), and (13) enforce limitations on the transmit power for each RU, ensure non-negative power allocation for each UE, and restrict the received power per UE, respectively.
Constraint (14) guarantees the QoS by ensuring that each service meets its minimum required data rate. Furthermore, (15) and (16) account for the fronthaul capacity limitations of each RU and the delay constraints for each UE, respectively, by their QoS requirements.
Constraints (17) and (18) define the binary nature of the decision variables $\alpha$ and $\beta$. Finally, constraints (19), (20), and (21) ensure that each UE is associated with exactly one RU and that Resource Block (RB) assignments are only made if the UE is associated with the corresponding RU while preventing multiple UEs from sharing the same RB on the same RU.
The system model is constrained by limited resources, such as UE or RU, and slice power and fronthaul capacity, which prevents the aggregate throughput from exceeding its optimal value. This ensures that the objective function, representing aggregate throughput, has an upper bound and cannot increase infinitely. As a result, the algorithm used to solve the system will converge to an optimal solution. If the objective function is a strictly increasing function concerning the number of iterations, it will converge to its global optimum. The algorithm will converge to a local optimum if the function is non-monotonically ascending. This guarantees the stability and convergence of the system model within the feasible region.
5 Methodology
To address the complex resource allocation problem in the O-RAN architecture, we introduce SS-VAE, a novel GAI framework rooted in semi-supervised learning. This method employs a specialized DNN architecture that intrinsically integrates a VAE —leveraging its ability to learn underlying data distributions—with a contrastive loss function designed to enhance representation quality. The SS-VAE is designed to comprehend complex relationships between input features and optimal resource allocation decisions (e.g., parameters $\alpha$, $\beta$, and transmission powers) that meet stringent QoS requirements. A comprehensive dataset underpins the development and initial supervised training of SS-VAE. This dataset features diverse O-RAN scenarios, with labeled data representing optimal allocation parameters generated via an Exhaustive Search Algorithm (ESA), which also serves as an optimal performance benchmark.